# Quantifying systematic uncertainties in white dwarf cooling age determinations

### v1

@@ 2.2 Parameter space constraints
From these fits, we defined the parameter space for our core oxygen profile. For each of the eight Akima spline parameters, we determined the minimum and maximum values obtained from fitting this diverse set of models and asteroseismological inferences. We then defined a uniform distribution spanning from this minimum to maximum for each parameter. This approach allows us to capture the range of predictions from current evolutionary calculations and asteroseismological analyses, reasonably encompassing the uncertainties in core oxygen profiles. The ranges of values explored for each parameter are given in Table 1.333While we have not explicitly included MESA models in the building of these distributions, we verified that they encompass the MESA models of Bauer (2023). For simplicity, we use the same range of values for the core oxygen parameters across all white dwarf masses. The meaning of each parameter is defined in Giammichele et al. 2017 (see their Figure 1); as examples, represents the size of the inner homogeneous core (in terms of ) and is the oxygen abundance in that same region.
@@
For the envelope composition, we vary the helium and hydrogen layer masses within ranges again informed by evolutionary calculations and observational constraints (Renedo et al., 2010; Romero et al., 2013; Giammichele et al., 2022). The upper limits for these envelope masses are set to ensure the model remains a white dwarf rather than a nuclear-burning AGB star, with specific limits depending on the white dwarf mass (Table 1).
@@ 2.3 Constitutive physics
The final uncertainty we consider relates to 22Ne shell distillation, a process which may significantly delay white dwarf cooling but is not included in standard evolutionary models. To account for this uncertainty while maintaining compatibility with existing cooling models, we adopt a two-component approach. First, we calculate a baseline systematic uncertainty using our ensemble of models without distillation, capturing the spread due to variations in core composition, envelope masses, and conductive opacities. To quantify the potential impact of shell distillation, we create a second ensemble by applying a cooling delay to all of our original models. The magnitude of this delay is drawn from a uniform distribution between 0.5 and 2 Gyr for each model, consistent with theoretical predictions (Segretain, 1996; Blouin et al., 2020a, 2021). The effective temperature range in which this delay is introduced is determined using both theoretical considerations and observational evidence. Shell distillation is expected to occur after the white dwarf has begun crystallizing but before crystallization is complete. Blouin et al. (2021) suggested that shell distillation takes place when % of the core is crystallized assuming a uniform and core composition. To account for uncertainties in core composition, we assume that the additional delay is injected between the effective temperatures corresponding to 35% and 65% crystallization on average in our cooling models. This aligns well with observational evidence from the mass– diagram presented by Kilic et al. (2020), which shows an overdensity of white dwarfs in certain temperature ranges at high mass (Figure 4). This feature cannot be fully explained by standard crystallization cooling delays and may provide evidence for 22Ne distillation.
@@ 3.2 Uncertainties with distillation
To interpret Table 2, it is useful to consider a few examples. For a 0.8 white dwarf at K, where no distillation can possibly have taken place because core crystallization has yet to begin, we report a simple systematic uncertainty of 2.6%. However, at 6000 K and below, distillation could have introduced an additional cooling delay. In these cases, we report two values: the first represents the baseline systematic uncertainty, while the value in parentheses indicates the potential additional delay due to distillation. The amplitude of this delay initially increases sharply with decreasing , reflecting the fact that the range in which we suspect shell distillation to be taking place is quite narrow. In relative terms, this shift then slowly decreases at lower temperatures, since the cooling age continues to increase while the distillation delay remains constant.
@@ acknowledgments
The MESA computations presented here were performed on the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto. The data analysis was carried out on the Astrohub virtual research environment (https://astrohub.uvic.ca/) hosted on the Digital Research Alliance of Canada Arbutus Cloud at the University of Victoria.

### v2

@@ 2.2 Parameter space constraints
From these fits, we defined the parameter space for our core oxygen profile. For each of the eight Akima spline parameters, we determined the minimum and maximum values obtained from fitting this diverse set of models and asteroseismological inferences. We then defined a uniform distribution spanning from this minimum to maximum for each parameter. This approach allows us to capture the range of predictions from current evolutionary calculations and asteroseismological analyses, reasonably encompassing the uncertainties in core oxygen profiles. The ranges of values explored for each parameter are given in Table 1.333While we have not explicitly included MESA models in the building of these distributions, we verified that they encompass the MESA models of Bauer (2023). For simplicity, we use the same range of values for the core oxygen parameters across all white dwarf masses. This approach is required due to the fragmentary nature of available constraints, as published models and asteroseismology data exist only for specific masses and a handful of stars. The meaning of each parameter is defined in Giammichele et al. 2017 (see their Figure 1); as examples, represents the size of the inner homogeneous core (in terms of ) and is the oxygen abundance in that same region.
@@
For the envelope composition, we vary the helium and hydrogen layer masses within ranges again informed by evolutionary calculations and observational constraints (Renedo et al., 2010; Romero et al., 2013; Giammichele et al., 2022). For both these parameters, the true probability distributions remain unknown, and we therefore adopt uniform distributions for simplicity. The upper limits for these envelope masses are set to ensure the model does not undergo nuclear burning (as expected for the vast majority of white dwarfs, Althaus et al., 2015), with specific limits depending on the white dwarf mass (Table 1).
@@ 2.3 Constitutive physics
The final uncertainty we consider relates to 22Ne shell distillation, a process which may significantly delay white dwarf cooling but is not included in standard evolutionary models. To account for this uncertainty while maintaining compatibility with existing cooling models, we adopt a two-component approach. First, we calculate a baseline systematic uncertainty using our ensemble of models without distillation, capturing the spread due to variations in core composition, envelope masses, and conductive opacities. To quantify the potential impact of shell distillation, we create a second ensemble by applying a cooling delay to all of our original models. The magnitude of this delay is drawn from a uniform distribution between 0.5 and 2 Gyr for each model, consistent with theoretical predictions (Segretain, 1996; Blouin et al., 2020a, 2021). The effective temperature range in which this delay is introduced is determined using both theoretical considerations and observational evidence. Shell distillation is expected to occur after the white dwarf has begun crystallizing but before crystallization is complete. Blouin et al. (2021) suggested that shell distillation takes place when % of the core is crystallized assuming a uniform and core composition. To account for uncertainties in core composition (which determines the at which shell distillation is expected to start, Blouin et al., 2021), we assume that the additional delay is injected between the effective temperatures corresponding to 35% and 65% crystallization on average in our cooling models. This is consistent with the observational mass– diagram presented by Kilic et al. (2024), which shows an overdensity of white dwarfs in certain temperature ranges at high mass. This feature cannot be fully explained by standard crystallization cooling delays and may provide evidence for 22Ne distillation.
@@ 3.1 Uncertainties without distillation
We also performed an additional analysis where we excluded the asteroseismological results from Giammichele et al. (2022) when defining the parameter space for the core oxygen profile. Using only the theoretical stellar evolution models to inform our parameter distributions yielded results very similar to those shown in Figure 5 (see Figure A1). Our uncertainty estimates are therefore not significantly impacted by the inclusion or exclusion of asteroseismological constraints on the core composition, which is consistent with our finding that the core composition is not the dominant source of uncertainty in white dwarf cooling ages.
@@ 3.2 Uncertainties with distillation
To interpret Table 2, it is useful to consider a few examples. For a 0.8 white dwarf at K, where no distillation can possibly have taken place because core crystallization has yet to begin, we report a simple systematic uncertainty of 2.6%. However, at 6000 K and below, distillation could have introduced an additional cooling delay. In these cases, we report two values: the first represents the baseline systematic uncertainty, while the value in parentheses indicates the potential additional delay due to distillation. The amplitude of this delay initially increases sharply with decreasing , reflecting the fact that the range in which we suspect shell distillation to be taking place is quite narrow. In relative terms, this shift then slowly decreases at lower temperatures, since the cooling age continues to increase while the distillation delay remains constant. The corresponding absolute uncertainties in Gyr are provided in Table A1.
@@ acknowledgments
We thank the anonymous reviewer for useful comments. Claude 3.5 Sonnet and ChatGPT were used to improve wording at the sentence level and assist with coding. The MESA computations presented here were performed on the Niagara supercomputer at the SciNet HPC Consortium. SciNet is funded by Innovation, Science and Economic Development Canada; the Digital Research Alliance of Canada; the Ontario Research Fund: Research Excellence; and the University of Toronto. The data analysis was carried out on the Astrohub virtual research environment (https://astrohub.uvic.ca/) hosted on the Digital Research Alliance of Canada Arbutus Cloud at the University of Victoria.